Perform each of the following conversions: (Must show all steps) (ABCDh (86257)s = ( 6. a. Simplify the following (Must show all steps) A+ BC+D (E+F) z= X Y + W Y + wX+XYZ(Reduce to Two terms)
Answer #1

ABCDis in Hexadecimal number system and wantto translate it into BCD. ABCD16 10.163+11.16*+12-161+13.16-40960+2816+192+13-43986257, in 9-ary number system and want to 86257,-8ダ+6ダ+2.945 947f = 52488+4374+162+45+7 = 570762, Happened: 57076 Converting